• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Bo ( RC )

    Log entry time 08:51:15 on October 03, 2009

    Entry number 292875

    keyword=Experiment update (8:45am 10/03/09)

    Beam was down since 2:00am this morning due to a RF problem. After that,
    around 3:00am, another problem, water leak, was found, so accelerator
    dropped from beam permit to power permit due to this power supply safety
    problem. MCC was trying to fix it. I called Sasha to cancel 9:00am Moller
    measurement. I will talk to Kent to schedule another one.

    A copy of this log entry has been emailed to: paschke, glamazdi, saha